# Stripe dark solitons in a nonlinear optical fiber with second- and fourth-order dispersion

### 11 Dec 2023 v1

@@ III SDS
Another problem worthy of consideration is the transmission stability of SDSs. From the insets of Fig. 2, one can see that some spontaneous oscillations emerge in the propagating process of SDSs, which is a widespread phenomenon in the nonlinear systems with MI. It prompts us to analyze the MI in the model (2). As we know, different types of perturbations can produce different phenomena under the influence of MI. Generally speaking, a purely periodic perturbation can generate Akhmediev breathers Dudley-2014 ; Dudley-2009 ; a purely localized one can generate a rogue wave Zhao-2016-JOSAB or Kuznetsov-Ma breather Duan-2019 , and induce the spontaneous oscillations in a triangular region El-1993 ; Biondini-2016 . (The word "purely" in front of "perturbation" is used to distinguish them with a periodic-localized perturbation.) A SDS at the initial distance can be seen as a periodic-localized perturbation on a plane wave, so we need to separately consider the influence of MI on its periodic and localized components. For its periodic component, we know that the traditional growing rate $g$ can be obtained by the modified growing rate $G$ when $\eta=0$, namely
$g(\omega)$ $=G(\omega,\eta=0)$
$=\omega^{2}(\frac{\beta_{2}}{2}+\frac{\beta_{4}}{24}\omega^{2})(2a_{0}^{2}+\frac{\beta_{2}}{2}\omega^{2}+\frac{\beta_{4}}{24}\omega^{4}).$ (17)
@@
If the SDS’s oscillation frequency $\omega_{s}$ locates in the MI band, its periodicity will impact the stability of perturbation, whose corresponding MI distribution is shown as the upper plot in Fig. 3 (c). Thus, we calculate the growing rate when $\omega=\omega_{s}$ by
$g(\omega_{s})=\frac{(3\beta_{2}^{2}-4a_{0}^{2}\beta_{4})(-27\beta_{2}^{2}+24\sqrt{3}a_{0}\beta_{2}\sqrt{\beta_{4}}+20a_{0}^{2}\beta_{4})}{64\beta_{4}^{2}},$ (18)
@@
to describe the MI growing rate induced by the periodicity. Because it can produce MI phenomena faster than the localization component, we call the periodicity-induced MI "strong MI", and accordingly we call the localization-induced MI "weak MI". The typical distribution of growing rate of weak MI is shown as the lower plot in Fig. 3 (c). In Fig. 3 (d), we illustrate the phase diagram of strong MI, weak MI, and modulation stability (MS) on the $\beta_{2}$-$\beta_{4}$ plane, and the growing rate of strong MI is presented by the color scale. The strong MI corresponds to the region of $\beta_{2}<0$ and $\frac{3\beta_{2}^{2}}{4a_{0}^{2}}<\beta_{4}<\frac{27\beta_{2}^{2}}{4a_{0}^{2}}$, and the weak MI is in the region of $\beta_{2}<0$ and $\beta_{4}>\frac{27\beta_{2}^{2}}{4a_{0}^{2}}$. When $\beta_{4}=1$, the strong and weak MI respectively need the conditions of $-2/\sqrt{3}<\beta_{2}<-2/9$ and $\beta_{2}>-2/9$. It is consistent with the evolution plots in Fig. 2, where the SDS is more and more stable with $\beta_{2}$ increasing.
@@
Besides of the fundamental structures shown above, we also find some SDSs with compound structures. When $\beta_{2}=0$, $\beta_{4}=1$, and $a_{0}=1$, the distributions of wave functions of SDSs with compound structures are shown as the upper row in Fig. 4. There are many kinds of compound structures, which have different number of humps, like (a) two small humps, (b) six small humps, (c) two big humps, (d) six big humps, and (e) two small and two big humps. The compound structures of SDS include but are not limited to these five kinds. Their counterparts without a phase step also exist but are not shown in this paper. Next, we study their numerical evolution to test their transmission stability, which are shown as the lower row in Fig. 4. The SDSs with compound structures can always propagate for a long distance, and no MI phenomenon emerges.
@@ IV SBS
As we know, the bright soliton with stripe structures in the model (2) has been studied in Refs. Zakharov-1998 ; Tam-2020 , including their existence condition and oscillation frequency. Here, we will study them by the MLSA method to provide a more accurate prediction for their characteristics and our understanding of their formation mechanism.
Figure 5: (Color online) (a) Distribution of the SBS’s wave function $\psi$ (red dashed curves) and their amplitude $|\psi|$ (blue solid curves), when $\beta_{2}=0$, $\beta_{4}=-1$ and $\beta=1$. (b) Same as (a) except for $\beta_{2}=0.75$. (c) In the $\omega$-$\eta$ plane, the characteristic parameters $(\omega_{s},\eta_{s})$ is the intersection point between the curves of $V_{\rm en}=0$ (black solid curve) and different $\beta$ value. The blue solid, purple dashed, and red solid curves denote $\beta=1$, $3/8$, and $1/8$, respectively. The parameters are $\beta_{2}=0.5$ and $\beta_{4}=-1$. (d) Dependence of SBS’s oscillation frequency $\omega_{s}$ on $\beta_{2}$, when $\beta_{4}=1$ and $\beta=1$. The red dots, the black solid curve, and the dashed blue curve are respectively the results of numerical solutions, the MLSA method [see Eq. (IV)], and the multiple scale analysis giving the expression $\omega_{s}=\sqrt{6\beta_{2}/|\beta_{4}|}$.
Firstly, by the biconjugate gradient method, we show the distribution of wave functions of SBS in Fig. 5 (a) when $\beta_{2}=0$, $\beta_{4}=-1$, and the propagation constant $\beta=1$. A bright soliton with oscillating tails can be seen, which is just the so-called pure-quartic soliton. Then, when we change $\beta_{2}$ to $0.75$, a bright soliton with many stripes appear. Unlike the SDS, the SBS do not have the background of plane wave. Thus, our calculation of MLSA directly starts with the perturbation,
@@
It is worth noting that the formation mechanism of stripes in SDS and SBS is different. For SDS, the stripes originate from the interaction between a perturbation and a zero-velocity plane wave; for SBS, the stripes are caused by the interaction between two components with opposite frequencies. Thus, the velocity of stripes in SBS is not equal to $V_{\rm st}=K/\omega$. On the other hand, the formation of localized envelopes is not related to the background wave, so $V_{\rm en}$ can still describe the velocity of localized envelopes in SBS. Its expression is
@@
$\omega_{s}={\rm Re}[\sqrt{\frac{3\beta_{2}+\sqrt{-6\beta\beta_{4}}}{-\beta_{4}}}],$
$\eta_{s}={\rm Re}[\sqrt{\frac{-3\beta_{2}+\sqrt{-6\beta\beta_{4}}}{-\beta_{4}}}].$ (24)
As shown in Fig. 5, the point $(\omega_{s},\eta_{s})$ is the intersection of the black curve ($V_{\rm en}=0$) and the blue curve ($\beta=1$), when $\beta_{2}=0.5$ and $\beta_{4}=-1$. We also depict the curves of $\beta=3/8$ and $\beta=1/8$: the former is the critical case where the intersection exists just right, and the latter has no intersection with the curve of $V_{\rm en}=0$. It indicates that the SBS cannot exist when $\beta<3/8$, and we also obtain the same result in our numerical simulations. Then, considering that $\omega_{s}\neq 0$ and $\eta_{s}\neq 0$, we can obtain that the existence condition of SBS is
@@
as shown in Fig. 1. We also show the oscillation frequency of SDS under different sets of $\beta_{2}$ in Fig. 5 (d), when $\beta_{4}=1$ and $\beta=1$. There are the results from three approaches, i.e., the numerical solutions, the MLSA method [see Eq. (IV)], and the multiple scale analysis in Ref. Tam-2020 giving the expression $\omega_{s}=\sqrt{6\beta_{2}/|\beta_{4}|}$. The MLSA method shows a better agreement with numerical results than the multiple scale analysis.
Figure 6: (Color online) (a) Dependence of $V_{\rm en}$ on $\beta_{4}$ when $\omega=2.1678$, $\eta=0.4466$, and $\beta_{2}=0.75$. The insets are the amplitude evolution of the initial state when $\beta_{4}=-2$, $-1$, and $-0.5$. (b) Dependence of $V_{\rm en}$ on $\omega$ when $\eta=0.4466$, $\beta_{2}=0.75$, and $\beta_{4}=-1$. The insets are the amplitude evolution of the initial state when $\omega=1$, $2.1678$, and $3$.
Table 1: Characteristics of different nonlinear waves
@@
The wave function (19) has two components with opposite frequencies. When both of their localized envelopes have the zero velocity, i.e., $V_{\rm en}=0$, a SBS will be generated, which can be considered as the result of synchronous transmission of the two components. To confirm our guess, for a wave function with frequency $\omega$ and steepness $\eta$, we show the dependence of $V_{\rm en}$ on $\beta_{4}$ in Fig. 6 (a), when $\omega=2.1678$, $\eta=0.4466$, and $\beta_{2}=0.75$. When $\beta_{4}=-1$, we have $V_{\rm en}=0$. The initial state has the form of
@@
Finally, we summary the characteristics of different nonlinear waves in Tab. 1. They contains the traditional solitons, the stripe solitons, the pure-quartic solitons, and periodic waves. We recall that the quantity $\varphi$ is defined by $\varphi=\tan^{-1}({\omega_{s}}/{\eta_{s}})$ to describe the strength of a wave’s periodicity relative to its localization, and $\omega_{s}$ and $\eta_{s}$ are the parameters determining the periodicity and localization of a wave, respectively. A traditional soliton has the localization but no the periodicity (i.e., $\omega_{s}=0$ and $\eta_{s}>0$), while a periodic wave has the periodicity but no the localization (i.e., $\omega_{s}>0$ and $\eta_{s}=0$). A stripe soliton has both of the periodicity and localization, and therefore it corresponds to $\omega_{s}>0$ and $\eta_{s}>0$. The pure-quartic soliton is a particular case of the stripe solitons when the periodicity and localization achieve a balance, i.e., $\omega_{s}=\eta_{s}$, so its proportion of periodicity and localization is $\omega_{s}/\eta_{s}=1$ and the characteristic quantity is $\varphi=\pi/4$.
@@ V Conclusion
In summary, the SDSs in nonlinear fibers with the second-order dispersion and FOD are investigated. The SDSs have the fundamental structures, their counterpart without a phase step, and some compound structures with different numbers of humps. With the help of the MLSA method, we quantitatively analyze the periodicity and localization of fundamental SDSs under different coefficients of second-order dispersion and FOD. Their oscillation frequency is successfully predicted and the corresponding MI is analyzed. For the SBSs, though their major characteristics have been studied in the previous works Zakharov-1998 ; Tam-2020 , we provide a new perspective to understand their formation mechanism and show a more accurate prediction of their oscillation frequency. Our method and results can provide the theoretical guidance for the generations and manipulations of stripe solitons in nonlinear fibers, and furthermore bring more possibility for their applications in soliton communications and optical measurements. Considering that stripe solitons have been also found in spinor Bose-Einstein condensates Achilleos-2013 ; Xu-2013 ; Zhao-2020 ; Yang-2021 , we expect that our method can be applied in these systems and help with the observation of matter-wave stripe solitons.
@@ Acknowledgement
The authors thank Prof. Jie Liu for their helpful discussions. This work was supported by NSAF (No.U2330401) and National Natural Science Foundation of China (No. 12247110).

### 07 Mar 2024 v2

@@ III.3 Analysis of propagation stability
Another problem worthy of consideration is the propagation stability of SDSs. From the insets of Fig. 2, one can see that some spontaneous oscillations emerge in the propagating process of SDSs. It is a widespread phenomenon in the nonlinear systems with MI, and therefore prompts us to analyze the MI in the model (2). As we know, different types of perturbations can produce different phenomena under the influence of MI. Generally speaking, a purely periodic perturbation can generate Akhmediev breathers Dudley-2014 ; Dudley-2009 ; a purely localized one can generate a rogue wave Zhao-2016-JOSAB or Kuznetsov-Ma breather Duan-2019 , and induce the spontaneous oscillations in a triangular region El-1993 ; Biondini-2016 . (The word "purely" in front of "perturbation" is used to distinguish them with a periodic-localized perturbation.) A SDS at the initial distance can be seen as a periodic-localized perturbation on a plane wave, so we need to separately consider the influence of MI on its periodic and localized components. For its periodic component, we know that the traditional growing rate $g$ can be obtained by the modified growing rate $G$ when $\eta=0$, namely
$g(\omega)$ $=|G(\omega,\eta=0)|$
$=\Big{|}\omega\;{\rm Re}\Big{[}\sqrt{-(\frac{\beta_{2}}{2}+\frac{\beta_{4}}{24}\omega^{2})(2\beta+\frac{\beta_{2}}{2}\omega^{2}+\frac{\beta_{4}}{24}\omega^{4}})\Big{]}\Big{|}.$ (25)
@@
If the SDS’s oscillation frequency $\omega_{s}$ locates in the MI band, its periodicity will intensively impact the stability of plane wave, whose corresponding MI distribution is shown as the upper plot in Fig. 3 (c). Thus, we calculate the growing rate when $\omega=\omega_{s}$ by
$g(\omega_{s})$ $=\frac{1}{8\beta_{4}}\sqrt{-\sqrt{3}\beta_{2}+2\sqrt{\beta\beta_{4}}}$
×$\times\sqrt{\sqrt{3}\beta_{2}(-9\beta_{2}^{2}+68\beta\beta_{4})+2\sqrt{\beta\beta_{4}}(9\beta_{2}^{2}+20\beta\beta_{4})},$ (26)
@@
to describe the MI growing rate induced by the periodicity of SDSs. Because it can produce MI phenomena faster than the localization component, we call the periodicity-induced MI "strong MI", and accordingly we call the localization-induced MI "weak MI". The typical distribution of growing rate of weak MI is shown as the lower plot in Fig. 3 (c). In Fig. 3 (d), we illustrate the phase diagram of strong MI, weak MI, and modulation stability (MS) on the $\beta_{2}$-$\beta_{4}$ plane, and the growing rate of strong MI is presented by the color scale. The strong MI corresponds to the region of $\beta_{2}<0$ and $\frac{3\beta_{2}^{2}}{4\beta}<\beta_{4}<\frac{27\beta_{2}^{2}}{4\beta}$, and the weak MI is in the region of $\beta_{2}<0$ and $\beta_{4}>\frac{27\beta_{2}^{2}}{4\beta}$. When $\beta_{4}=1$, the strong and weak MI respectively need the conditions of $-2/\sqrt{3}<\beta_{2}<-2/9$ and $\beta_{2}>-2/9$. It is consistent with the evolution plots in Fig. 2, where the SDS is more and more stable with $\beta_{2}$ increasing.
III.4 Multi-soliton bound state
@@
In this subsection, we focus on the multi-dark-soliton bound states. In Ref. Buryak-1995 , the bound states of several bright solitons under the FOD have been found, and their existence condition and stability were analyzed. Here, we will find the bound states of dark solitons by the biconjugate gradient method Press-book . When $\beta_{2}=0$, $\beta_{4}=1$, and $\beta=1$, the distributions of wave functions of the multi-dark-soliton bound states are shown as the upper row in Fig. 4. There are many kinds of bound states, which have different number of amplitude humps, such as (a) two small humps, (b) six small humps, (c) two big humps, (d) six big humps, and (e) two small and two big humps. The bound states include but are not limited to these five kinds, and their counterparts without the total phase step also exist but are not shown in this paper. Next, we study their numerical evolution to test their propagation stability, which are shown as the lower row in Fig. 4. These bound states can always propagate for a long distance, and no MI phenomenon emerges.
We stress that the “dispersion waves” in Figs. 2 (a-c) are the manifestation of the MI of plane-wave backgrounds, which is induced by the unstable frequency sideband from the periodicity of dark solitons. As shown in Fig. 3 (a), the MI can be strong, weak, or absent in different cases, and therefore the dark solitons can propagate for a short or long distance under the influence of MI. Considering that they could be observed in the optical fibers with different length, our calculations about dark solitons can provide the theoretical guidance for their observations. Particularly, when $\beta_{4}\geq 0$, the dark solitons and their bound states are stable and propagate for a long distance, as shown in Figs. 2 (d,e) and 4. In addition, the traditional linear stability analysis can also be used to obtain the linear equation (5) and the classical dispersion law (19). However, they can give us only the MI condition of a perturbed plane wave, but cannot give anything about striped solitons. This is the reason why we use the MLSA method to analyze the SDSs.
Figure 5: (Color online) (a) Distribution of the SBS’s wave function $\psi$ (red dashed curves) and its amplitude $|\psi|$ (blue solid curves), when $\beta_{2}=0$, $\beta_{4}=-1$ and $\beta=1$. (b) Same as (a) except for $\beta_{2}=0.75$. (c) In the $\omega$-$\eta$ plane, the characteristic parameters $(\omega_{s},\eta_{s})$ is the intersection point between the curves of $V_{\rm en}=0$ (black solid curve) and different $\beta$ value. The blue solid, purple dashed, and red solid curves denote $\beta=1$, $3/8$, and $1/8$, respectively. The parameters are $\beta_{2}=0.5$ and $\beta_{4}=-1$. (d) Dependence of SBS’s oscillation frequency $\omega_{s}$ on $\beta_{2}$, when $\beta_{4}=1$ and $\beta=1$. The red dots and the black solid curve are respectively the results of numerical solutions and the MLSA method [see Eq. (32)].
@@ IV SBS
As we know, the bright soliton with stripe structures in the model (2) has been studied in Refs. Zakharov-1998 ; Tam-2020 , including their existence condition and oscillation frequency. Here, we will study them by the MLSA method and then provide a possible interpretation of their formation mechanism.
Firstly, by the biconjugate gradient method, we show the distribution of wave functions of SBS in Fig. 5 (a) when $\beta_{2}=0$, $\beta_{4}=-1$, where the propagation constant is $\beta=1$. A bright soliton with oscillating tails can be seen, which is just the so-called pure-quartic soliton Blanco-2016 . Then, when we change $\beta_{2}$ to $0.75$, a bright soliton with many stripes appears in Fig. 5 (b). Unlike the SDS, the SBS has on background of plane wave. Thus, our calculation of MLSA directly starts with the perturbation,
@@
It is worth noting that the formation mechanism of stripes in SDS and SBS is different. For SDS, the stripes originate from the interaction between a perturbation and a zero-velocity plane wave; for SBS, the stripes are caused by the interaction between two components with opposite frequencies. Thus, the velocity of stripes in SBS is not equal to $V_{\rm st}=K/\omega$. Nevertheless, the formation of localized envelopes is not related to the background wave, so $V_{\rm en}$ can still describe the velocity of localized envelopes in SBS. Its expression is
@@
$\omega_{s}={\rm Re}[\sqrt{\frac{3\beta_{2}+\sqrt{-6\beta\beta_{4}}}{-\beta_{4}}}],\;\eta_{s}={\rm Re}[\sqrt{\frac{-3\beta_{2}+\sqrt{-6\beta\beta_{4}}}{-\beta_{4}}}].$ (32)
As shown in Fig. 5 (c), the point $(\omega_{s},\eta_{s})$ is the intersection of the black curve ($V_{\rm en}=0$) and the blue curve ($\beta=1$), when $\beta_{2}=0.5$ and $\beta_{4}=-1$. We also depict the curves of $\beta=3/8$ and $\beta=1/8$: the former is the critical case where the intersection exists just right, and the latter has no intersection with the curve of $V_{\rm en}=0$. It indicates that the SBS cannot exist when $\beta<3/8$, and we also obtain the same result in our numerical simulations. Then, considering that $\omega_{s}\neq 0$ and $\eta_{s}\neq 0$, we can obtain that the existence condition of SBS is
@@
as shown in Fig. 1. We also show the oscillation frequency of SDS under different sets of $\beta_{2}$ in Fig. 5 (d), when $\beta_{4}=1$ and $\beta=1$. Our analytical prediction (32) shows a good agreement with the numerical result.
Figure 6: (Color online) (a) Dependence of $V_{\rm en}$ on $\beta_{4}$ when $\omega=2.1678$, $\eta=0.4466$, and $\beta_{2}=0.75$. The insets are the amplitude evolution of the initial state (34) when $\beta_{4}=-2$, $-1$, and $-0.5$. (b) Dependence of $V_{\rm en}$ on $\omega$ when $\eta=0.4466$, $\beta_{2}=0.75$, and $\beta_{4}=-1$. The insets are the amplitude evolution of the initial state (34) when $\omega=1$, $2.1678$, and $3$.
Table 1: Characteristics of several nonlinear waves
@@
Our derived existence condition (33) and oscillation frequency (32) of SBSs agree well with the results given by Ref. Tam-2020 . It indicates that for SBSs, the MLSA method are equivalent to the direct asymptotic analysis, due to the absence of plane-wave background. Nevertheless, besides of the above results, the MLSA method can also provide the interpretation for the formation mechanism of SBSs. It is known that the wave function (27) has two components with opposite frequencies. When both of their localized envelopes have the zero velocity, i.e., $V_{\rm en}=0$, a SBS will be generated. It means that the formation of a SBS is the result of synchronous propagation of the two components. To confirm this guess, for a wave function with frequency $\omega$ and steepness $\eta$, we show the dependence of $V_{\rm en}$ on $\beta_{4}$ in Fig. 6 (a), when $\omega=2.1678$, $\eta=0.4466$, and $\beta_{2}=0.75$. When $\beta_{4}=-1$, we have $V_{\rm en}=0$. The initial state has the form of
@@
Finally, we summarize the characteristics of different nonlinear waves in Tab. 1. They contains the traditional solitons, the striped solitons, the pure-quartic solitons, and periodic waves. We recall that the quantity $\varphi$ is defined by $\varphi=\tan^{-1}({\omega_{s}}/{\eta_{s}})$ to describe the strength of a wave’s periodicity relative to its localization, and $\omega_{s}$ and $\eta_{s}$ are the parameters determining the periodicity and localization of a wave, respectively. A traditional soliton has the localization but no the periodicity (i.e., $\omega_{s}=0$ and $\eta_{s}>0$), while a periodic wave has the periodicity but no the localization (i.e., $\omega_{s}>0$ and $\eta_{s}=0$). A striped soliton has both of the periodicity and localization, and therefore it corresponds to $\omega_{s}>0$ and $\eta_{s}>0$. The pure-quartic soliton is a particular case of the striped solitons when the periodicity and localization achieve a balance, i.e., $\omega_{s}=\eta_{s}$, so its proportion of periodicity and localization is $\omega_{s}/\eta_{s}=1$ and the characteristic quantity is $\varphi=\pi/4$.
@@ V Conclusion
In summary, the excitations of dark solitons in a nonlinear fiber with the second-order dispersion and FOD are investigated. In this system, we find the SDSs, their counterparts without the total phase step, and some multi-dark-soliton bound states. With the help of the MLSA method, we quantitatively analyze the periodicity and localization of the SDSs under different coefficients of second-order dispersion and FOD. Their oscillation frequency is successfully predicted and the corresponding MI is analyzed, which can make the propagation of SDSs unstable. For the SBSs, though their major characteristics have been studied in the previous works Zakharov-1998 ; Tam-2020 , we provide a new possible perspective to understand their formation mechanism. Our method and results can provide the theoretical guidance for the generations and manipulations of striped solitons in nonlinear fibers, and furthermore bring more possibility for their applications in soliton communications and optical measurements. Considering that stripe solitons have been also found in spinor Bose-Einstein condensates Achilleos-2013 ; Xu-2013 ; Zhao-2020 ; Yang-2021 , we expect that our method can be applied in these systems and help with the observation of matter-wave striped solitons. We also note that the competition between periodicity and localization of a soliton is similar to the well-known wave-particle duality to some extent. Understanding the similarity between them is a challenging but might be important task, which is worthy of future’s study.
@@ Acknowledgement
The authors thank Prof. Jie Liu for his helpful discussion. This work was supported by NSAF (No.U2330401) and National Natural Science Foundation of China (No. 12247110).